OpenStudy (anonymous):

A rectangle's length is 4 feet more than its width. write a quadratic function that expresses the rectangle's area in terms of its width.

OpenStudy (solomonzelman):

length=l width=w l=w+4 normally the Area of a triangle is height * Base *1/2 assuming length is height, and base is width l*w *(1/2) substitute w+4 for l (w+4)w(1/2) 

OpenStudy (anonymous):

Answer choices are: A(w)=w+4 A(w)=w^2+4w A(w)=lw A(w)=w^2-4w I'm guessing its the first.

OpenStudy (solomonzelman):

 but since it is are of a triangle it is 1/2 of that. maybe you meant for area of rectangle, not triangle?